# TriangleMix: A Lossless and Efficient Attention Pattern for Long Context Prefilling

### v1

@@ 3.3 Analysis of Start Layer
Ours + FP (0.95) 3.4 (-17%) 5.2 (-29%) 7.2 (-36%) 9.2 (-42%) 11.3 (-47%) 13.6 (-51%) 16.1 (-53%)
FlexPrefill (0.90) 3.3 (-20%) 5.1 (-30%) 6.8 (-39%) 8.9 (-44%) 11.1 (-48%) 13.6 (-51%) 15.5 (-55%)
Ours + FP (0.90) 3.3 (-20%) 4.5 (-38%) 6.7 (-40%) 8.5 (-47%) 10.5 (-51%) 12.6 (-54%) 14.5 (-58%)
Table 6: Time-to-first-token (TTFT) in seconds measured on Llama-3.1-8B-Instruct. We separate FlexPrefill ($\gamma=0.90$) as it generally leads to accuracy degradation, whereas the other settings do not.
@@ 3.4 Efficiency of TriangleMix
We implement Triangle Attention using Triton Tillet et al. (2019), with further implementation details provided in Appendix A.2. All experiments are conducted on a single NVIDIA A100 80GB GPU. Dense attention is implemented using FlashAttention Kwon et al. (2023).
Attention Kernel Latency. We first benchmark the average attention kernel time per layer on Llama-3.1-8B-Instruct at sequence lengths of 32K, 64K, and 128K. As shown in Table 5, Triangle Attention significantly outperforms dense attention, achieving 3.7×3.7$\times$3.7 × to 15.3×15.3$\times$15.3 × speedup. The speedup mainly comes from Triangle Attention’s linear complexity ($O(N)$), which makes it much more efficient than $O(N^{2})$ dense attention. Besides, unlike dynamic sparsity methods, it doesn’t require estimating attention block indices during runtime, which saves additional overhead.
End-to-End TTFT. We further measure time-to-first-token (TTFT) across sequence lengths of 32K, 48K, 64K, 96K, 112K, and 128K. Results for Llama-3.1-8B-Instruct are reported in Table 6. TriangleMix applies dense attention in shallow layers and Triangle Attention in deeper layers, reducing TTFT by 12%–32%. It is also compatible with dynamic attention for enhanced performance. For example, integrating MInference with TriangleMix reduces TTFT from 18.0s to 14.5s (a 19% decrease). Under the lossless setting, TriangleMix combined with FlexPrefill ($\gamma=0.95$) achieves the fastest TTFT for 32K–80K sequences, while TriangleMix with MInference is optimal for 96K–128K. We visualize the results of some selected methods in Figure 6.
4 Discussion
In this section, we discuss why Middle Q-K attention can be safely removed. Our hypothesis is that the observed phenomena arise from a Train–Test Misalignment. During training, the loss is applied uniformly to all positions in the input sequence. For example, if the input contains 2048 tokens, the model is also trained to predict token 1001 given tokens 1–1000. However, during inference, the primary goal is to predict tokens after the prompt (such as token 2049), making predictions of tokens inside the prompt (such as token 1001) largely irrelevant.
To validate this assumption, we conduct an experiment on Llama-3.1-8B-Instruct. We sample 100 Wikipedia articles and extract the first 2048 tokens of each. In the last 16 layers, we apply Triangle Attention, skipping attention for Q indices between 512 and 1024. We then measure changes of the average perplexity (PPL) relative to dense attention in two output spans.
Method PPL (1024–1152) PPL (1920–2048)
Dense 8.02 7.79
TriangleMix 8.34 (+0.32) 7.81 (+0.02)
Table 7: With 2048 input tokens, we measure perplexity on two output spans. Triangle attention is applied, skipping attention between Q indices 512–1024 in the last 16 layers. Experiments are on Llama-3.1-8B-Instruct.
The results in Table 7 show that skipping attention of Q index 512–1024 in deep layers noticeably increases perplexity when predicting tokens in the 1024–1152 range. For these positions, the skipped attention effectively becomes the Last Q-K attention necessary for predicting the next token. In contrast, for positions beyond 1920, the effect is minimal.
The Middle Q-K region is important primarily for predicting tokens within the prompt, while for generation tasks focused on tokens after the prompt, the Middle Q-K in deeper layers is largely redundant. Although further theoretical work is needed, this observation suggests that selective sparsification of attention during inference can be done safely.
@@ 5 Related Works
Static Sparsity Attention. Early methods employ fixed sparse patterns, such as strided Child et al. (2019), dilated Ding et al. (2023), sliding window Jiang et al. (2023), and mixed patterns Beltagy et al. (2020), typically requiring training from scratch. DuoAttention Xiao et al. (2024) introduces head-level static sparsity but needs additional offline training. Training-free approaches like Streaming Xiao et al. (2023) and Triangle Li et al. (2024a) attention often degrade accuracy for long contexts. In contrast, the proposed TriangleMix attention pattern significantly mitigates performance loss and is nearly lossless in accuracy.
Dynamic Sparsity Attention. Methods like MInference Jiang et al. (2024a), FlexPrefill Lai et al. (2025), and XAttention Xu et al. (2025) dynamically determine attention masks online, maintaining high accuracy but incurring computational overhead for moderately long contexts. TriangleMix can enhance these dynamic approaches for further acceleration .
Long-context LLM Inference. FlashAttention Dao et al. (2022) speeds up attention by reducing memory access through fused operations. PagedAttention Kwon et al. (2023) improves decoding by managing KV cache allocation efficiently. KV cache optimizations techniques also include token-level eviction (SnapKV Li et al. (2024b)) and query-aware cache selection (Quest Tang et al. (2024)). TriangleMix is orthogonal to these approaches.
@@ 6 Conclusion
In this paper, we propose a hybrid attention pattern for long-context LLMs, where dense attention is used in shallow layers and Triangle attention is applied in deeper layers. This pattern is training-free and achieves nearly lossless accuracy. Our method achieves a 12% to 32% time-to-first-token reduction across 32K to 128K inputs. Moreover, it can be further accelerated when combined with dynamic sparsity attention.
Limitations
One limitation of our method is that the sparsity ratio is fixed within the pretrained model. We observe that Llama-3.1-8B-Instruct and Llama-3-8B-Instruct-262K exhibit higher inherent sparsity compared to Qwen2.5-7B-Instruct, which constrains the achievable acceleration on the latter. It may be possible to enhance the sparsity of models like Qwen2.5-7B-Instruct through a post-training optimization process, which we leave for future exploration.
@@ Appendix A Appendix
A.1 Task to Probe Attention Importance
We use a key-value retrieval task to analyze the importance of different attention sections. The language model is prompted to retrieve the correct value, beginning from its first generated token. The prompt is as follows:
@@ A.1 Task to Probe Attention Importance
We randomly generate pairs of keys and values using Version 4 UUID strings.
@@ A.2 Implementation of Triangle Attention
We implement Triangle Attention using Triton Tillet et al. (2019), with details provided in Algorithm 1. For each row index, we apply a different attention mechanism: if the row index is less than $N-N_{\text{last}}$, we apply Streaming Attention; otherwise, we apply chunked FlashAttention to increase GPU utilization. The outputs from both segments are then merged to produce the final result.
@@
 Initialize $\boldsymbol{O}\leftarrow(0)^{N\times d_{h}}$
 Initialize output buffer $\boldsymbol{O}_{\mathrm{last}}\leftarrow(0)^{SN_{\mathrm{last}}\times d_{h}}$
 Initialize LSE buffer $\boldsymbol{l}_{\mathrm{last}}\leftarrow(-\inf)^{SN_{\mathrm{last}}}$
@@
  Initialize $\boldsymbol{O}_{\text{chip}}\leftarrow(0)^{B_{M}\times d_{h}}$
  Initialize $\boldsymbol{m}\leftarrow(-\inf)^{B_{N}}$
  Initialize $\boldsymbol{s}\leftarrow(0)^{B_{M}}$
@@
    Load $\boldsymbol{K}_{\text{chip}}\leftarrow\boldsymbol{K}^{jB_{N}:(j+1)B_{N}}$
    Load $\boldsymbol{V}_{\text{chip}}\leftarrow\boldsymbol{V}^{jB_{N}:(j+1)B_{N}}$
@@
    Load $\boldsymbol{K}_{\text{chip}}\leftarrow\boldsymbol{K}^{jB_{N}:(j+1)B_{N}}\in\mathbb{R}^{B\times d_{h}}$
    Load $\boldsymbol{V}_{\text{chip}}\leftarrow\boldsymbol{V}^{jB_{N}:(j+1)B_{N}}\in\mathbb{R}^{B\times d_{h}}$
@@
    Load $\boldsymbol{K}_{\text{chip}}\leftarrow\boldsymbol{K}^{jB_{N}:(j+1)B_{N}}$
    Load $\boldsymbol{V}_{\text{chip}}\leftarrow\boldsymbol{V}^{jB_{N}:(j+1)B_{N}}$
@@
A.3 Additional Longbench Data
Table 8 compares our method with dynamic sparsity attention on Longbench. When combined with MInference, our method achieves similar performance to MInference alone. Integrating with FlexPrefill yields a slight improvement over its original results.
Methods Average
2Wiki
GovRep
Hotpot
LCC
MNews
MF-en
PsgCnt
PsgRtr
Qasper
Rbench
Samsum
TREC
TrvQA
[…]
FlexPrefill; $\gamma=0.90$ 44.0 31.0 32.8 48.8 59.1 25.1 51.3 4.0 39.0 37.1 50.8 41.5 65.3 86.2
[…]
FlexPrefill; $\gamma=0.95$ 48.4 39.4 33.7 58.3 67.2 25.6 55.6 4.0 47.7 42.5 53.8 41.9 69.7 90.2
[…]
FlexPrefill; $\gamma=0.90$ 33.6 16.8 32.7 23.0 46.8 25.6 45.6 2.3 8.0 23.2 41.4 36.8 61.0 73.3
[…]
FlexPrefill; $\gamma=0.95$ 35.0 17.7 33.0 25.1 36.8 26.1 49.1 3.3 13.7 28.3 36.2 39.1 65.0 81.7
[…]
FlexPrefill; $\gamma=0.90$ 37.3 17.5 29.3 27.6 64.0 21.8 39.0 7.0 28.7 15.1 51.4 39.3 64.7 79.3
[…]
FlexPrefill; $\gamma=0.95$ 40.8 20.2 29.7 33.9 66.8 21.8 40.4 4.0 48.7 17.1 54.8 40.5 67.3 85.5
[…]
Table 8: Comparison with dynamic sparsity methods on LongBench.

### v2

@@ 3.3 Analysis of $L_{\mathrm{tri}}$
FlexPrefill 3.6 (-12%) 5.5 (-25%) 7.7 (-31%) 9.9 (-38%) 12.3 (-42%) 15.0 (-45%) 17.8 (-48%)
Ours + FlexPrefill 3.4 (-17%) 5.2 (-29%) 7.2 (-36%) 9.2 (-42%) 11.3 (-47%) 13.6 (-51%) 16.1 (-53%)
XAttention 4.1 (-1%) 6.6 (-10%) 9.4 (-16%) 12.4 (-22%) 15.7 (-26%) 19.3 (-30%) 23.2 (-33%)
Ours + XAttention 3.5 (-14%) 5.5 (-24%) 7.7 (-31%) 10.0 (-37%) 12.4 (-42%) 15.1 (-45%) 17.8 (-48%)
Table 7: Time-to-first-token (TTFT) in seconds measured on Llama-3.1-8B-Instruct.
@@ 3.4 Efficiency of TriangleMix
We implement Triangle Attention using Triton (Tillet et al., 2019), with further implementation details provided in Appendix A.2. All experiments are conducted on a single NVIDIA A100 80GB GPU. Dense attention is implemented using FlashAttention (Kwon et al., 2023).
Attention Kernel Latency. We first benchmark the average attention kernel time per layer on Llama-3.1-8B-Instruct and Qwen2.5-7B-Instruct at sequence lengths of 32K, 64K, and 128K. As shown in Table 6, Triangle Attention achieves $3.7\times$ to $15.3\times$ speedup compared to dense attention,. The speedup mainly comes from Triangle Attention’s linear complexity ($O(N)$), which makes it much more efficient than $O(N^{2})$ dense attention. Besides, unlike dynamic sparsity methods, TriangleMix does not require estimating attention block indices. Its triangle attention kernel is also much simpler to implement, which further reduces overhead.
End-to-End TTFT. We measure the end-to-end time-to-first-token (TTFT) for sequence lengths of 32K, 48K, 64K, 96K, 112K, and 128K. Table 7 reports the results on Llama-3.1-8B-Instruct. TriangleMix yields a TTFT reduction of 12%–32%. Compared with DuoAttention, TriangleMix achieves similar efficiency but causes less performance degradation (see Section 3.2). In addition, DuoAttention partitions heads within the same layer into different sparsity patterns (Liu et al., 2025), which leads to imbalance under tensor parallelism, while TriangleMix avoids this issue. Moreover, TriangleMix can be seamlessly combined with dynamic attention for further efficiency gains. For instance, integrating MInference with TriangleMix lowers TTFT from 18.0s to 14.5s (a 19% reduction) at length 128K. TriangleMix combined with FlexPrefill achieves the lowest TTFT for 32K–80K inputs, whereas TriangleMix with MInference is optimal for 96K–128K inputs.
@@ 4 Related Works
Static Sparsity Attention. Early methods employ fixed sparse patterns, such as strided (Child et al., 2019), dilated (Ding et al., 2023), sliding window (Jiang et al., 2023), and mixed patterns (Beltagy et al., 2020), typically requiring training from scratch. DuoAttention (Xiao et al., 2024) introduces head-level static sparsity but needs additional offline training. Streaming pattern is a training-free approach (Xiao et al., 2023) but leads to degraded accuracy for long contexts (Li et al., 2024a). In contrast, the proposed TriangleMix attention pattern significantly mitigates performance loss and is nearly lossless in accuracy.
Dynamic Sparsity Attention. Existing methods such as MInference (Jiang et al., 2024a), FlexPrefill (Lai et al., 2025), and XAttention (Xu et al., 2025) dynamically identify attention blocks with high scores and restrict computation to these blocks. While this form of attention score sparsity focuses on pruning blocks with small attention scores, our work moves beyond by uncovering a different kind of sparsity, which we term decoding-time contribution sparsity. Specifically, we show that many blocks with non-trivial scores still contribute little to decoding, and thus can be safely skipped. Leveraging this property enables additional acceleration in the prefilling stage.
Long-context LLM Inference. FlashAttention (Dao et al., 2022) speeds up attention by reducing memory access through fused operations. PagedAttention (Kwon et al., 2023) improves decoding by managing KV cache allocation efficiently. KV cache optimizations techniques also include token-level eviction (SnapKV (Li et al., 2024b)) and query-aware cache selection (Quest (Tang et al., 2024)). TriangleMix is orthogonal to these approaches.
@@ 5 Conclusion
In this paper, we identify a new form of sparsity in the prefilling stage, termed decoding-time contribution sparsity, and introduce TriangleMix, a training-free static attention pattern. By selectively applying Triangle attention in certain layers, TriangleMix substantially reduces attention overhead while maintaining nearly lossless performance. On 128K inputs, it achieves up to a 15.3× speedup in attention computation, surpassing typical dynamic sparse methods. Moreover, TriangleMix can be seamlessly combined with dynamic sparsity, yielding an additional 6%–19% reduction in TTFT.
@@ Appendix A Appendix
A.1 Details of Attention Block Probing
We divide the causal attention matrix into three distinct sections as illustrated in Figure 1(a):
  • Streaming section: includes attention sink and the sliding window;
  • Last Q–K section: covers interactions between the last part of $\bm{Q}$ and $\bm{K}$, excluding the Streaming section;
  • Middle Q–K section: consists of the remaining interactions between the middle parts of $\bm{Q}$ and $\bm{K}$.
We define the Streaming mask as:
where $si$ is the number of sink tokens and $sl$ is the sliding window size.
We define the Last Q–K mask as:
where $\textit{last}\geq 1$ specifies the number of rows corresponding to the last section.
The Middle Q–K mask is defined as:
In our gradient-based attention block probing experiments, we set the parameters to $si=64$, $sl=128$, and $last=128$. We adopt a key-value retrieval task, where the language model is prompted to output the correct value starting from its very first generated token. Each input sequence contains approximately 2,000 tokens, and gradients are computed over 100 randomly generated samples. The prompt used in this task is as follows:
@@ A.1 Details of Attention Block Probing
We randomly generate pairs of keys and values using UUID strings.
@@ A.2 Implementation of Triangle Attention
We implement Triangle Attention using Triton (Tillet et al., 2019), with details provided in Algorithm 1. For each row index, we apply a different attention mechanism: if the row index is less than $N-N_{\text{last}}$, we apply Streaming Attention; otherwise, we apply chunked FlashAttention to increase GPU utilization. The outputs from both segments are then merged to produce the final result.
@@
 Initialize $\boldsymbol{O}\leftarrow(0)^{N\times d_{h}}$, output buffer $\boldsymbol{O}_{\mathrm{last}}\leftarrow(0)^{SN_{\mathrm{last}}\times d_{h}}$, LSE buffer $\boldsymbol{l}_{\mathrm{last}}\leftarrow(-\inf)^{SN_{\mathrm{last}}}$
@@
  Initialize $\boldsymbol{O}_{\text{chip}}\leftarrow(0)^{B_{M}\times d_{h}}$, $\boldsymbol{m}\leftarrow(-\inf)^{B_{N}}$, $\boldsymbol{s}\leftarrow(0)^{B_{M}}$
@@
    Load $\boldsymbol{K}_{\text{chip}}\leftarrow\boldsymbol{K}^{jB_{N}:(j+1)B_{N}}$, $\boldsymbol{V}_{\text{chip}}\leftarrow\boldsymbol{V}^{jB_{N}:(j+1)B_{N}}$
@@
    Load $\boldsymbol{K}_{\text{chip}}\leftarrow\boldsymbol{K}^{jB_{N}:(j+1)B_{N}}\in\mathbb{R}^{B\times d_{h}}$, $\boldsymbol{V}_{\text{chip}}\leftarrow\boldsymbol{V}^{jB_{N}:(j+1)B_{N}}\in\mathbb{R}^{B\times d_{h}}$
@@
    Load $\boldsymbol{K}_{\text{chip}}\leftarrow\boldsymbol{K}^{jB_{N}:(j+1)B_{N}}$, $\boldsymbol{V}_{\text{chip}}\leftarrow\boldsymbol{V}^{jB_{N}:(j+1)B_{N}}$
@@
A.3 Acknowledgement of LLM Usage
We used large language models to polish the writing of this paper, and all generated content was carefully reviewed to ensure precise expression.